OpenStudy (anonymous):

Given log 10 = 1 and log 2 = a what can we afirm? a) log 40 = 1 + 2^a b) log 40 = 2 + a c) log 5 = a - 1 d) log 5 = 2 - a So I proved wrong both a and b: log 40 => log 4 + log 10 = 2log 2 + 1 => 2a + 1 I don't know what to do with the other options. Help pls!

myininaya (myininaya):

\[\log(40)=\log(2^2 \cdot 10)=\log(2^2)+\log(10)=2\log(2)+\log(10)=2a+1\] \[\log(5)=\log(\frac{10}{2})=\log(10)-\log(2)=1-a\]

myininaya (myininaya):

I'm not getting any of that

OpenStudy (bahrom7893):

i was gonna say that too.. i sat here solving that nonsense for a while..

myininaya (myininaya):

lol poor bahrom

OpenStudy (anonymous):

Me too. Weird... Thx guys

OpenStudy (anonymous):

i don't get it :(